The perimeter of a rectangle is 82 centimeters. The length is 3 cm more than the width. Find the dimensions of the rectangle.
Answer by myonlinemathhelpdotcom(3)

The perimeter of a rectangle is 82 centimeters. The length is 3 cm more than the width. Find the dimensions of the rectangle.
2*l + 2*w = 82
l=3+w
2*(3+w) + 2*w = 82
6+2*w + 2*w = 82
6 + 4*w = 82
4*w = 76
w = 19
l = 3+w = 22
my online math help dot com